OpenStudy (anonymous):

Find the square. (7m - 5)^2 How do you solve one of these problems?

OpenStudy (anonymous):

49m^2 - 70m +25

OpenStudy (anonymous):

(a+b)^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

Thanks a lot.

OpenStudy (anonymous):

(a-b)^2 = a^2 - 2ab + b^2